OpenStudy (anonymous):

Solve the quadratic equation by factorisation 2x^2 + 5 x + 2 = 0 Any help will be appreciated.

OpenStudy (anonymous):

(2x +1)(x + 2)

OpenStudy (amistre64):

do you know why factoring is easier to solve it with?

OpenStudy (anonymous):

So is (2x + 1) (x+ 2) = 0 2x + 1 = 0 2x = -1 / 2 = -0.5 The other x+2 = 0 x = 2 Would that be right?

OpenStudy (amistre64):

thats good, but x = -2 for the second one ... you mighta had a typo

OpenStudy (anonymous):

Many Thanks amiste I just needed confirmation that I was using the right method!